Given f(x) = 3x + 2, find f (a – 1). Please HELP!!!
Answer by elima(1433)

Given f(x) = 3x + 2, find f (a – 1)
f(a-1)=3(a-1)+2
f(a-1)=3a-3+2
f(a-1)=3a-1
:)